# ViMoE: An Empirical Study of Designing Vision Mixture-of-Experts

### v1

@@ 6 Conclusion
In this work, we integrate the sparse Mixture-of-Experts (MoE) architecture into the classic Vision Transformer (ViT), termed ViMoE, to explore its potential application in image classification. We report the challenges encountered in designing ViMoE, particularly in determining the configuration of MoE layers without prior guidance, as inappropriate expert arrangements can negatively impact convergence. To mitigate this, we introduce the shared expert to stabilize the training process, thus streamlining the design by eliminating the need for repeated trials to find the optimal configuration. Furthermore, by observing the routing behavior and the distribution of samples across experts, we identify the MoE layers that are crucial for the divide-and-conquer processing of data. These insights allow us to refine the ViMoE architecture, achieving both efficiency and competitive performance. We hope this work provides new insights into the design of MoE models for vision tasks and offers valuable empirical guidance for future research.

### v2

@@ 7 Conclusion
In this work, we integrate the sparse Mixture-of-Experts (MoE) architecture into the classic Vision Transformer (ViT), termed ViMoE, to explore its potential application in computer vision tasks. We report the challenges encountered in designing ViMoE, particularly in determining the configuration of MoE layers without prior guidance, as inappropriate expert arrangements can negatively impact convergence. To mitigate this, we introduce the shared expert to stabilize the training process, thus streamlining the design by eliminating the need for repeated trials to find the optimal configuration. Furthermore, by observing the routing behavior and the distribution of samples across experts, we identify the MoE layers crucial for handling data in a divide-and-conquer manner. These insights allow us to refine the ViMoE architecture, achieving both efficiency and competitive performance. We hope this work provides new insights into the design of MoE models for vision tasks and offers valuable empirical guidance for future research.